# Improved High-probability Convergence Guarantees of Decentralized SGD

### v1

@@ 3.1 Preliminaries
Assumption (A1) is satisfied by connected undirected graphs, as well as a class of strongly-connected directed graphs with doubly stochastic weights, e.g., [24]. Moreover, it can be shown that (A1) implies that the second largest singular value of $W$, denoted by $\lambda$, satisfies $\lambda\in[0,1)$, see [72].
@@
Assumptions (A2)-(A3) are standard in smooth non-convex optimization, e.g., [29, 11, 25].
@@
The first condition in (A4) is standard in decentralized stochastic optimization [25, 53], while the second and third require noise to be unbiased and light-tailed (i.e., sub-Gaussian). Light tails are necessary to achieve exponentially decaying tail bounds of the form (1) if the vanilla stochastic gradient estimator is employed without any modifications (e.g., using clipping operator or estimators like median-of-means) and are widely used in centralized settings, e.g., [28, 29, 30, 39, 31, 33, 32]. [^6] While it is possible to achieve concentration of the form in (1) under, e.g., sub-Weibull noise, see [41] and references therein, the focus of this work is on the more widely studied sub-Gaussian noise setting.
@@
A heterogeneity bound of the type (A5) is in fact required to ensure that DSGD converges for non-convex costs [73, 8, 11, 26]. Note that (A5) is strictly weaker than the uniformly bounded gradient assumption used in [57], as it allows users’ gradients to grow with the global gradient. Compared to [11], who analyze the MSE convergence and require a bound on the average heterogeneity, i.e., $\frac{1}{n}\sum_{i\in[n]}\|\nabla f_{i}(x)\|^{2}\leq A^{2}+B^{2}\|\nabla f(x)\|^{2}$, we impose a slightly stronger condition. While we believe that (A5) can be relaxed to average heterogeneity by incorporating a similar analysis technique to the one in [11], to keep the proofs simple and instructive, we will use assumption (A5).
@@
Assumption (A6) is used in the analysis of strongly convex costs. It is well known that DSGD requires (A3) and (A6) to hold for each $f_{i}$, e.g., [20, 11, 26] which is in general not required for centralized SGD. The heterogeneity bound in (A5) and strong convexity of each cost in (A6) can be removed by deploying the GT technique, e.g., [25], however, this is beyond the scope of the current work.
@@
Lemma 2 establishes some properties of noise, importantly showing that the average noise is $\mathcal{O}\Big(\frac{\sigma\sqrt{d}}{\sqrt{n}}\Big)$-sub-Gaussian, highlighting the variance reduction benefit of decentralized optimization in the HP sense. This result is crucial toward showing that DSGD achieves a linear speed-up in the number of users, which we establish in the following subsections. We note that the dependence on problem dimension $d$ is unavoidable due to the definition of sub-Gaussianity used in (A4), with further discussion provided in Section 4 and the Appendix.
@@ 3.2 Non-convex Costs
Let (A1)-(A5) hold. If for any $T\geq 1$, the step-size is chosen such that $\alpha_{T}\equiv\alpha=\min\left\{C,\frac{\sqrt{n}}{\sigma\sqrt{dT}}\right\}$, where $C>0$ is a problem related constant satisfying
@@
$\frac{1}{nT}\sum_{t\in[T]}\sum_{i\in[n]}\|\nabla f(x^{t}_{i})\|^{2}$ $=\mathcal{O}\Bigg(\frac{\sigma\sqrt{d}\big(\Delta_{f}+\log(\nicefrac{{1}}{{\delta}})+L\big)}{\sqrt{nT}}+\frac{L^{2}\Delta_{x}}{(1-\lambda^{2})T}$
$+\frac{\Delta_{f}+\log(\nicefrac{{1}}{{\delta}})}{CT}+\frac{n\lambda^{2}(A^{2}+\sigma^{2})]}{\sigma^{2}(1-\lambda^{2})T}\Bigg).$
Theorem 1 establishes HP convergence guarantees for DSGD with general smooth non-convex costs and a fixed step-size. The bound in Theorem 1 consists of three terms, with the first two matching centralized SGD, while the third captures the effect of the network, through $\frac{\lambda^{2}}{1-\lambda^{2}}$. Next, note that the leading term $\mathcal{O}\big(\frac{\sigma\sqrt{d}}{\sqrt{nT}}\big)$ achieves a linear speed-up in the number of users, showing the benefits of decentralized learning in the HP sense. Finally, we can see that, in the case $A=0$, the above bound recovers the rates of noiseless centralized gradient descent (GD), i.e., when $\sigma=0$, the above bound becomes $\mathcal{O}\big(\frac{1}{T}\big)$. Compared to [57], where the authors provide a $\mathcal{O}\big(\frac{\log(\nicefrac{{T}}{{\delta}})}{\sqrt{T}}\big)$ HP convergence bound, while requiring uniformly bounded gradients, we provide several improvements, by relaxing the uniformly bounded gradients assumption, achieving a linear speed-up and recovering the rates of GD when $\sigma=0$ and the gradients of users grow at most as the global gradient.
@@ 3.3 Strongly Convex Costs
In this subsection we provide improved results for the last iterate of strongly convex costs. It is well known that strong convexity implies a unique global minimizer of $f$, see, e.g., [74]. Denote the global minimizer by $x^{\star}\in\mathbb{R}^{d}$, noting that $f^{\star}\coloneqq f(x^{\star})$ and let $\kappa\coloneqq\frac{L}{\mu}\geq 1$ and $\|\nabla\mathbf{f}^{\star}\|^{2}\coloneqq\sum_{i\in[n]}\|\nabla f_{i}(x^{\star})\|^{2}$ be the condition number and heterogeneity measure, respectively. Next, recall that DSGD does not require a heterogeneity bound to converge in the MSE sense for strongly convex costs, e.g., [20, 11]. To show similar benefits in the HP sense, a different approach to Lemma 3 is required. This is achieved by carefully bounding the MGF of the consensus gap and leveraging properties of strong convexity. The formal result is stated next.
@@
Lemma 4 gives a tight bound on the MGF of the consensus gap, without requiring bounded heterogeneity. Using Lemma 6 from the Appendix, it can be shown that the RHS is of the order $e^{\mathcal{O}(\nu K_{t+1}\alpha_{t}^{2})}$. We note that the requirement of same model initialization across users can be removed, at the cost of an additional term in the exponent on the RHS, of order $\mathcal{O}(\nu K_{t+1}\lambda^{t}\Delta_{x})$, which decays geometrically (recall that $\lambda\in[0,1)$). However, we make the assumption for ease of exposition. Prior to stating the main theorem, we provide an important technical result, which allows us to establish sharper bounds and ensure linear speed-up in the number of users is achieved.
@@
Lemma 5 provides a bound on the MGF of an “almost decreasing” process. Compared to bounds used in the centralized setting, e.g., [31, 32, 40], Lemma 5 gives a more fine-grained result, allowing us to incorporate higher-order terms in the final rate. We are now ready to state the main result for strongly convex costs.
@@
Theorem 2 establishes HP convergence guarantees for smooth strongly convex costs and time-varying step-size, without requiring the bounded heterogeneity condition (A5). In the Appendix we provide the full bound, containing additional terms, of order $\mathcal{O}\big(\frac{\log(T+t_{0}))}{(T+t_{0})^{3}}\big)$. We can see that the leading term in Theorem 2 decays at a rate $\mathcal{O}\big(\frac{d\sigma^{2}\kappa}{n(T+t_{0})}\big)$, achieving linear speed-up, while the network, problem and heterogeneity dependence are captured through the higher-order term $\mathcal{O}\big(\frac{\lambda^{2}\kappa^{2}\|\nabla\mathbf{f}^{\star}\|^{2}}{(1-\lambda)^{2}n(T+t_{0})^{2}}\big)$. Crucially, the term without linear speed-up, i.e., $\mathcal{O}\big(\frac{\sigma^{2}}{(T+t_{0})^{2}}\big)$, achieves a faster decay, stemming from the improved result in Lemma 5. Compared to [57], where the authors establish the $\mathcal{O}\big(\frac{1}{T}\big)$ convergence rate in HP for PL costs, Theorem 2 provides several improvements. First, in [57] the authors require uniformly path-wise bounded gradients, i.e., $\|\nabla f_{i}(x^{t}_{i})\|\leq C_{i}$, for every $t\geq 1$ almost surely, which is difficult to guarantee uniformly across all sample paths. Secondly, they impose the following condition on the MGF of the noise
@@
where $\alpha_{t}=\frac{a}{t+t_{0}}$ is the step-size. Such condition implies that the noise is $\frac{\sigma_{i}}{t+t_{0}}$-sub-Gaussian at time $t$, meaning that the noise asymptotically vanishes, at rate $\mathcal{O}\big(\frac{1}{t+t_{0}}\big)$. [^8] It can be easily verified via Markov’s inequality that a $\frac{\sigma}{t+t_{0}}$-sub-Gaussian random variable $X$ satisfies $\mathbb{P}\big(X^{2}\leq\frac{\sigma^{2}(\log(\nicefrac{{1}}{{\delta}})+1)}{(t+t_{0})^{2}}\big)\geq 1-\delta$, for any $\delta\in(0,1)$. Similarly, it follows from Jensen’s inequality that $\mathbb{E}[X^{2}]\leq\frac{\sigma^{2}}{(t+t_{0})^{2}}$. We remove both requirements, showing that HP convergence of the last iterate of strongly convex functions is guaranteed without bounded gradients and under standard sub-Gaussian noise, while also achieving linear speed-up in the number of users. Finally, we believe that the techniques used in our proof can be leveraged to show HP guarantees for the broader class of PL costs, which is left for future work.
@@ 4 Proof Outlines and Discussion
We can use Lemma 3 to control the last term on the RHS. To deal with $\sum_{t\in[T]}\alpha_{t}\langle\nabla f(\overline{x}^{t}),\overline{z}^{t}\rangle$, while removing the need for uniformly bounded gradients, we use Lemma 2 and the “offset trick”, e.g., [39, 33, 46], by subtracting $\sum_{t\in[T]}\frac{9\sigma^{2}\alpha_{t}^{2}\|\nabla f(\overline{x}^{t})\|^{2}}{4n}$ from both sides of (4), ensuring that the effects of $\langle\nabla f(\overline{x}^{t}),\overline{z}^{t}\rangle$ are absorbed by the left-hand side, allowing us to show that the resulting MGF is bounded. The rest of the proof then relies on Lemma 2, some technical results, careful selection of the step-size and Hölder’s inequality, see the Appendix for details.
@@
Discussion. Results of Theorems 1 and 2 are strong, demonstrating that DSGD converges in the HP sense under the same conditions on the cost function as in the MSE sense, while retaining order-optimal rates and linear speed-up in the number of users. Compared to works studying HP convergence of centralized SGD, e.g., [31, 32, 33, 40], we make several contributions. First, we face the challenge of controlling the MGF of the consensus gap, stemming from the decentralized nature of the algorithm. To that end, we provide Lemma 3 for non-convex costs, relying on the bounded heterogeneity condition, and the improved Lemma 4 for strongly convex costs, which removes the bounded heterogeneity condition. Next, we provide Lemma 2, which shows the variance reduction benefits of decentralized methods in the HP sense, which is of independent interest for HP studies of decentralized methods. To ensure linear speed-up in the number of users is achieved for strongly convex costs, we establish Lemma 5, which gives a more fine-grained result on the MGF [31, 32, 40], facilitating higher-order terms in the final bound, and is of independent interest. Moreover, to be able to utilize Lemma 5, a careful analysis of the MGF is needed, in order to simultaneously balance the effects of optimality and consensus gaps (recall that (5), which is used to bound the MGF of $F^{t+1}$, consists of both optimality and consensus gaps), while also ensuring that the “almost decreasing” property is maintained.
On the dimension dependence. We note that our bounds show a dependence on the problem dimension, of order $\sqrt{d}$ for non-convex and $d$ for strongly convex costs, which is not the case for either MSE or HP bounds, e.g., [11, 57]. This dependence stems from the third result in Lemma 2, where the problem dimension shows up in the sub-Gaussianity constant, which is unavoidable, due to the nature of the definition of sub-Gaussianity in (A4). In particular, to show the third property in Lemma 2, we use the fact that $\overline{z}^{t}$ is $\mathcal{O}\Big(\frac{\sigma\sqrt{d}}{\sqrt{n}}\Big)$-norm-sub-Gaussian, as a direct proof fails to yield the variance reduction benefit in the number of users. One way to avoid the (linear) dependence on $d$ is to assume that the noise is norm-sub-Gaussian, which would simultaneously relax the dependence on the dimension to $\log(d)$ and achieve the variance reduction benefit, at the cost of imposing a slightly stronger noise condition, see, e.g., Corollary 7 and the conclusion in [75]. [^10] See the Appendix for the full proof of Lemma 2 and further discussion on the dimension dependence.

### v2

@@ 3.1 Preliminaries
Assumption (A1) is satisfied by connected undirected graphs, as well as a class of strongly-connected directed graphs with doubly stochastic weights, e.g., [ran-vr]. Let $\lambda\coloneqq\|W-J\|$ denote the network connectivity parameter, where $J\coloneqq\frac{1}{n}\mathbf{1}_{n}\mathbf{1}_{n}^{\top}\in\mathbb{R}^{n\times n}$ is the ideal communication matrix. It can be shown that (A1) implies $\lambda\in[0,1)$, see [Horn_Johnson_2012].
@@
Assumptions (A2)-(A3) are standard in smooth non-convex optimization, e.g., [ghadimi2013stochastic, pmlr-v119-koloskova20a, ran-improved].
@@
The first condition in (A4) is standard in decentralized stochastic optimization [ran-improved, yu2025decentralized], while the second and third require noise to be unbiased and light-tailed (i.e., sub-Gaussian). Light tails are necessary to achieve exponentially decaying tail bounds of the form (1) if the vanilla stochastic gradient estimator is employed without any modifications (e.g., using clipping operator or estimators like median-of-means) and are widely used in centralized settings, e.g., [nemirovski2009robust, ghadimi2013stochastic, li-orabona, li2020high, pmlr-v99-harvey19a, liu2023high, pmlr-v206-bajovic23a]. [^7] While it is possible to achieve concentration of the form in (1) under, e.g., sub-Weibull noise, see [madden2020high] and references therein, the focus of this work is on the more widely studied sub-Gaussian noise setting.
@@
A heterogeneity bound of the type (A5) is in fact required to ensure that DSGD converges for non-convex costs [NIPS2017_f7552665, dist-learn-noncvx, pmlr-v119-koloskova20a, wang-cooperative]. Note that (A5) is strictly weaker than the uniformly bounded gradient assumption used in [dsgd-high-prob], as it allows users’ gradients to grow with the global gradient. Compared to [pmlr-v119-koloskova20a], who analyze the MSE convergence and require a bound on the average heterogeneity, i.e., $\frac{1}{n}\sum_{i\in[n]}\|\nabla f_{i}(x)\|^{2}\leq A^{2}+B^{2}\|\nabla f(x)\|^{2}$, we impose a slightly stronger condition. While we believe that (A5) can be relaxed to average heterogeneity by incorporating a similar analysis technique to the one in [pmlr-v119-koloskova20a], to keep the proofs simple and instructive, we will use assumption (A5).
@@
Assumption (A6) is used in the analysis of strongly convex costs. It is well known that DSGD requires (A3) and (A6) to hold for each $f_{i}$, e.g., [conv-rates-dsgd-jakovetic, pmlr-v119-koloskova20a, wang-cooperative] which is in general not required for centralized SGD. The heterogeneity bound in (A5) and strong convexity of each cost in (A6) can be removed by deploying the GT technique, e.g., [ran-improved], however, this is beyond the scope of the current work.
@@
Lemma 2 establishes some properties of noise, importantly showing that the average noise is $\mathcal{O}\Big(\frac{\sigma\sqrt{d}}{\sqrt{n}}\Big)$-sub-Gaussian, highlighting the variance reduction benefit of decentralized optimization in the HP sense. This result is crucial toward showing that DSGD achieves a linear speed-up in the number of users, which we do in the following subsections. We note that the dependence on problem dimension $d$ is unavoidable due to the definition of sub-Gaussianity used in (A4), with further discussion provided in Section 4 and the Appendix.
@@ 3.2 Non-convex Costs
Let (A1)-(A5) hold. If for any $T\geq 1$, the step-size is chosen such that $\alpha_{T}\equiv\alpha=\min\left\{C,\frac{\sqrt{n}}{\sigma\sqrt{dLT}}\right\}$, where $C>0$ is a problem related constant satisfying
@@
$\frac{1}{nT}$ $\sum_{t\in[T]}\sum_{i\in[n]}\|\nabla f(x^{t}_{i})\|^{2}=\mathcal{O}\Bigg(\frac{\sigma\sqrt{dL}\big(\Delta_{f}+\log(\nicefrac{{1}}{{\delta}})\big)}{\sqrt{nT}}$
$+\frac{\Delta_{f}+\log(\nicefrac{{1}}{{\delta}})}{CT}+\frac{L^{2}\Delta_{x}}{(1-\lambda^{2})T}+\frac{n\lambda^{2}L(A^{2}+\sigma^{2})}{\sigma^{2}d(1-\lambda)^{2}T}\Bigg).$
Theorem 1 establishes HP convergence guarantees for DSGD with general smooth non-convex costs and a fixed step-size. We now discuss the bound from a few perspectives.
Centralized SGD and linear speed-up. Compared to centralized SGD, which achieves the rate $\mathcal{O}\Big(\frac{1}{\sqrt{T}}+\frac{1}{T}\Big)$, e.g., [liu2023high], we can see that the rate from Theorem 1 similarly consists of two terms of different orders, namely $\mathcal{O}\Big(\frac{1}{\sqrt{nT}}+\frac{n}{T}\Big)$. Crucially, the leading term achieves a linear speed-up in the number of users, which is consistent with MSE results, e.g., [pmlr-v119-koloskova20a], and shows that DSGD maintains the benefits of decentralized learning, in the HP sense.
Network effect and transient time. We can see from the bound in Theorem 1 that network connectivity only affects the higher-order terms decaying at rate $\mathcal{O}\Big(\frac{1}{T}\Big)$, which is again consistent with MSE results, e.g., [pmlr-v119-koloskova20a]. Moreover, we can use the rate in Theorem 1 to establish the transient time of DSGD in the HP sense, which can be shown to be of the order $\mathcal{O}\Big(\frac{n^{3}}{(1-\lambda)^{4}}\Big)$, matching the transient time of DSGD stemming from MSE results, see, e.g., [unified-refined, Table I]. For details on the derivation of the transient time, see the Appendix.
Effects of noise and heterogeneity. The noise (through $\sigma,\overline{\sigma}$) and heterogeneity (through $A,B$) mainly affect the higher-order terms, [^9] The effect of noise on the leading term can be eliminated by choosing the step-size $\alpha=\min\Big\{C,\frac{\sqrt{n}}{\sqrt{dLT}}\Big\}$. which is again consistent with MSE results. Note that, if $A=0$, the bound in Theorem 1 recovers the convergence rate of noiseless gradient descent, in the sense that, when $\sigma=A=0$, the above bound becomes $\mathcal{O}\big(\frac{1}{T}\big)$.
Comparison with [dsgd-high-prob]. The authors in [dsgd-high-prob] study HP guarantees of DSGD over undirected time-varying networks, establishing a rate $\mathcal{O}\big(\frac{\log(\nicefrac{{T}}{{\delta}})}{\sqrt{T}}\big)$, using a time-varying step-size schedule $\alpha_{t}=\frac{C}{\sqrt{t+1}}$, for some $C>0$ and assuming uniformly bounded gradients. As we discuss next, our results provide several important improvements. First, we remove the uniformly bounded gradient assumption and replace it with the more general heterogeneity condition (A5), which is facilitated by the use of the “offset trick”, see the discussion in Section 4 for details. Next, the authors in [dsgd-high-prob] are unable to establish a linear speed-up, which is a byproduct of both their step-size choice, as well as their treatment of the noise. In particular, choosing a time-varying step-size for non-convex costs results in a loss of linear speed-up, as all terms converge at the same $\mathcal{O}\Big(\frac{1}{\sqrt{T}}\Big)$ rate. However, it would be impossible to obtain a linear speed-up in [dsgd-high-prob] even under a fixed step-size, as the authors show no variance reduction benefit of decentralized learning in the HP sense, instead only using the definition of sub-Gaussianity. On the other hand, the variance reduction benefits of decentralized learning established in Lemma 2 allow us to achieve linear speed-up in the number of users when a fixed step-size is used (see the Appendix for a detailed discussion on the results needed to achieve linear speed-up). [^10] This is also consistent with existing MSE guarantees, e.g., [pmlr-v119-koloskova20a, ran-improved, unified-refined], where linear speed-up for non-convex costs is achieved using a fixed step-size. Finally, while we consider static networks, our results can be readily extended to the time-varying setting considered in [dsgd-high-prob], see the Appendix for details.
@@ 3.3 Strongly Convex Costs
In this subsection we provide improved results for the last iterate of strongly convex costs. It is well known that strong convexity implies a unique global minimizer of $f$, see, e.g., [nesterov-lectures_on_cvxopt]. Denote the global minimizer by $x^{\star}\in\mathbb{R}^{d}$, noting that $f^{\star}\coloneqq f(x^{\star})$ and let $\kappa\coloneqq\frac{L}{\mu}\geq 1$ and $\|\nabla\mathbf{f}^{\star}\|^{2}\coloneqq\sum_{i\in[n]}\|\nabla f_{i}(x^{\star})\|^{2}$ be the condition number and heterogeneity measure, respectively. Next, recall that DSGD does not require a heterogeneity bound to converge in the MSE sense for strongly convex costs, e.g., [conv-rates-dsgd-jakovetic, pmlr-v119-koloskova20a]. To show similar benefits in the HP sense, a different approach to Lemma 3 is required. This is achieved by carefully bounding the MGF of the consensus gap and leveraging properties of strong convexity. The formal result is stated next.
@@
Lemma 4 gives a tight bound on the MGF of the consensus gap, without requiring the bounded heterogeneity condition (A5). Using Lemma 6 from the Appendix, it can be shown that the RHS is of the order $e^{\mathcal{O}(\nu K_{t+1}\alpha_{t}^{2})}$. We note that the requirement of same model initialization across users can be removed, at the cost of an additional term in the exponent on the RHS, of order $\mathcal{O}(\nu K_{t+1}\lambda^{t}\Delta_{x})$, which decays geometrically (recall that $\lambda\in[0,1)$). However, we assume same model initialization, for ease of exposition. Prior to stating the main theorem, we provide an important technical result, which allows us to establish sharper bounds and ensure linear speed-up in the number of users is achieved.
@@
Lemma 5 provides a bound on the MGF of an “almost decreasing” process. Compared to bounds used in the centralized setting, e.g., [pmlr-v99-harvey19a, pmlr-v206-bajovic23a, liu2024revisiting], Lemma 5 gives a more fine-grained result, allowing us to incorporate higher-order terms in the final rate, which is crucial for establishing linear speed-up, see Section 4 and the Appendix for detailed discussions. We are now ready to state the main result for strongly convex costs.
@@
Theorem 2 establishes HP convergence guarantees for smooth strongly convex costs and time-varying step-size, without requiring the bounded heterogeneity condition (A5). In the Appendix we provide the full bound, containing additional terms, of order $\mathcal{O}\big(\frac{\log(T+t_{0}))}{(T+t_{0})^{3}}\big)$. Next, we discuss multiple aspects of our results.
Centralized SGD and linear speed-up. Compared to centralized SGD, which achieves the HP rate $\mathcal{O}\Big(\frac{1}{T}+\frac{1}{T^{2}}\Big)$, e.g., [liu2024revisiting], the rate in Theorem 2 also consists of terms of different order, namely $\mathcal{O}\Big(\frac{1}{nT}+\frac{1}{T^{2}}\Big)$. Similarly to non-convex results, the leading term shows a linear speed-up in the number of users, again matching the corresponding MSE guarantees, e.g., [pmlr-v119-koloskova20a].
Network effect and transient time. We can see that network connectivity again only affects higher-order terms. Next, we can use the rate in Theorem 2 to establish a transient time of DSGD in the HP sense. As shown in the Appendix, our HP results imply a transient time of order $\mathcal{O}\Big(\max\Big\{\frac{n}{1-\lambda},\frac{1}{(1-\lambda)^{2}}\Big\}\Big)$, which is strictly sharper than the transient time of DSGD in, e.g., [unified-refined, Table II], stemming from MSE rates, which is of order $\mathcal{O}\Big(\frac{n}{(1-\lambda)^{2}}\Big)$. The sharper transient time further highlights the tightness of our results and an improved analysis.
Effects of noise and heterogeneity. While heterogeneity again affects only higher-order terms (through $\|\nabla\mathbf{f}^{\star}\|^{2}$ and $\|{\mathbf{x}}^{1}-{\mathbf{x}^{\star}}\|^{2}$, see the Appendix for the full bound), we can see that the noise affects both the leading and higher-order terms, stemming from the time-varying step-size schedule. Similarly to the discussion after Theorem 1, the effect of noise on the leading term can be mitigated by employing a carefully selected fixed step-size. Due to time-varying step-sizes being sufficient for achieving linear speed-up and for ease of exposition, the results using a fixed step-size are omitted.
Comparison with [dsgd-high-prob]. The authors in [dsgd-high-prob] establish HP convergence of DSGD for the more general case of PŁ costs, providing a rate of $\mathcal{O}\big(\frac{1}{T}\big)$, using the same time-varying step-size schedule considered in Theorem 2. Crucially, the authors require two very strong assumptions: $(i)$ path-wise uniformly bounded gradients, i.e., $\|\nabla f_{i}(x^{t}_{i})\|\leq G$, for all $t\geq 1$ and some $G>0$, almost surely; $(ii)$ asymptotically vanishing noise, i.e., for all $i\in[n]$ and any $t\geq 1$, the authors require the following condition
@@
where $\alpha_{t}=\frac{a}{t+t_{0}}$ is the step-size. Both conditions are very difficult to satisfy in practice, as $(i)$ is a path-wise relaxation of the uniformly bounded gradient condition (which is in general not satisfied even for strongly convex costs), while $(ii)$ implies that the noise is $\frac{\sigma_{i}}{t+t_{0}}$-sub-Gaussian at time $t$, meaning that the noise asymptotically vanishes, at rate $\mathcal{O}\big(\frac{1}{t+t_{0}}\big)$. [^11] It can be easily verified via Markov’s inequality that a $\frac{\sigma}{t+t_{0}}$-sub-Gaussian random variable $X$ satisfies $\mathbb{P}\Big(X^{2}\leq\frac{\sigma^{2}(\log(\nicefrac{{1}}{{\delta}})+1)}{(t+t_{0})^{2}}\Big)\geq 1-\delta$, for any $\delta\in(0,1)$. Similarly, it follows from Jensen’s inequality that $\mathbb{E}[X^{2}]\leq\frac{\sigma^{2}}{(t+t_{0})^{2}}$. Crucially, the authors in [dsgd-high-prob] are unable to establish linear speed-up in the number of users, even though much stronger assumptions on the noise and cost are imposed. On the other hand, our work uses standard assumptions, while simultaneously showing optimal convergence rates with linear speed-up and improved transient times compared to MSE results. These improvements stem not only from a sharper analysis and Lemma 2, but also Lemmas 4 and 5, which allow us to remove the bounded heterogeneity assumption (A5) and ensure linear speed-up is achieved for strongly convex costs (see the Appendix for a detailed discussion).
@@ 4 Proof Outlines and Discussion
We can use Lemma 3 to control the last term on the RHS. To deal with $\sum_{t\in[T]}\alpha_{t}\langle\nabla f(\overline{x}^{t}),\overline{z}^{t}\rangle$, while removing the need for uniformly bounded gradients, we use Lemma 2 and the “offset trick”, e.g., [li2020high, liu2023high, armacki2025optimal], by subtracting $\sum_{t\in[T]}\frac{9\sigma^{2}\alpha_{t}^{2}\|\nabla f(\overline{x}^{t})\|^{2}}{4n}$ from both sides of (4), ensuring that the effects of $\langle\nabla f(\overline{x}^{t}),\overline{z}^{t}\rangle$ are absorbed by the left-hand side, allowing us to show that the resulting MGF is bounded. The rest of the proof then relies on Lemma 2, some technical results, careful selection of the step-size and Hölder’s inequality, see the Appendix for details.
@@
Discussion. Results of Theorems 1 and 2 are strong, demonstrating that DSGD converges in the HP sense under the same conditions on the cost function as in the MSE sense, while retaining order-optimal rates and linear speed-up in the number of users. Compared to works studying HP convergence of centralized SGD, e.g., [pmlr-v99-harvey19a, pmlr-v206-bajovic23a, liu2023high, liu2024revisiting], we make several contributions. First, we face the challenge of controlling the MGF of the consensus gap, stemming from the decentralized nature of the algorithm. To that end, we provide Lemma 3 for non-convex costs, removing the uniformly bounded gradient condition and using the bounded heterogeneity condition (A5) instead, as well as the improved Lemma 4 for strongly convex costs, which removes (A5) altogether. Next, we provide Lemma 2, which shows the variance reduction benefits of decentralized methods in the HP sense and is of independent interest for HP studies of decentralized methods. To ensure linear speed-up in the number of users is achieved for strongly convex costs, we establish Lemma 5, which gives a more fine-grained result on the MGF of an “almost decreasing” process compared to existing (centralized) works, e.g., [pmlr-v99-harvey19a, pmlr-v206-bajovic23a, liu2024revisiting], facilitating higher-order terms in the final bound, and is of independent interest (see the Appendix for a detailed discussion on the importance of Lemma 5 for achieving linear speed-up for strongly convex costs). Finally, to be able to utilize Lemma 5, a careful analysis of the MGF is needed, in order to simultaneously balance the effects of optimality and consensus gaps (recall that (5), which is used to bound the MGF of $F^{t+1}$, consists of both optimality and consensus gaps), while also ensuring that the resulting process is “almost decreasing”, in the sense of Lemma 5.
On the dimension dependence. We note that our bounds show a dependence on the problem dimension, of order $\sqrt{d}$ for non-convex and $d$ for strongly convex costs, which is not the case for either MSE or HP bounds, e.g., [pmlr-v119-koloskova20a, dsgd-high-prob]. This dependence stems from the third result in Lemma 2, where the dimension figures in the sub-Gaussianity constant and is unavoidable, due to the definition of sub-Gaussianity used in (A4). In particular, to show the third property in Lemma 2, we use the fact that $\overline{z}^{t}$ is $\mathcal{O}\Big(\frac{\sigma\sqrt{d}}{\sqrt{n}}\Big)$-norm-sub-Gaussian, as a direct proof fails to yield the variance reduction benefit in the number of users. One way to avoid the (linear) dependence on $d$ is to assume that the noise is norm-sub-Gaussian, which would simultaneously relax the dependence on the dimension to $\log(d)$ and achieve the variance reduction benefit, at the cost of imposing a slightly stronger noise condition, see, e.g., Corollary 7 and the conclusion in [jin2019short]; see also the Appendix for a further discussion on the dimension dependence.